### v1

# Direct observation of the crystal electric-field splitting under magnetic field and uncovering field-induced magnetic phase transition in triangular rare-earth magnet CsErSe2

Abstract
An indispensable step toward understanding magnetic interaction in rare-earth magnets is the determination of spatially anisotropic single-ion properties resulting from the crystal electric field (CEF) physics. The CEF Hamiltonian exhibits a discrete energy spectrum governed by a set of independent parameters that reflect the site symmetry of the magnetic ion. However, experimentally determining these parameters for magnetic ions at low-symmetry sites has been proven highly challenging. In this study, we directly measured the CEF level splitting under magnetic fields ($\mathbf{B}$) using optical spectroscopy and extracted both CEF parameters and the exchange energies of a triangular insulating magnet CsErSe2 as a model system. With increasing field, we find many CEF levels undergo level-crossing which accompanies switching of the eigenstate. Particularly, such a crossing occurring at the ground state results in a step-like increase in magnetization that we captured with the low-temperature AC magnetic susceptibility measurements. Our work demonstrates that the accurately determined CEF Hamiltonian parameters enable uncovering the rich physics of field-induced collective magnetic phenomena, and potentially lead to a new route to the emergence of magnetic frustration.
I Introduction
Magnetic insulators with strong spin-orbit coupling are now widely recognized as a robust platform to investigate a range of quantum many-body phases, including quantum spin liquids (QSLs) and other phenomena, via various forms of magnetic frustration Savary and Balents (2016); Scheie et al. (2019).
In particular, compounds based on $4f$ rare-earth (RE) ions offer an excellent playground to explore various interesting magnetic interactions, due to their unique combination of geometric frustration arising from the two-dimensional triangular lattice and multiple magnetic degrees of freedom induced by spin-orbit coupling (SOC) and crystal electric field (CEF) interaction. Gardner et al. (2010); Rau and Gingras (2019); Bordelon et al. (2019); Scheie et al. (2024); Paddison et al. (2017); Gao et al. (2018); Li et al. (2015); Mauws et al. (2018); Scheie et al. (2020).
In RE-based insulating magnets, however, detailed understanding of the microscopic physics is a prerequisite for developing effective spin models that serve as a basis for theories of many-body phenomena. For a free RE ion, the spin ($S$) and orbital ($L$) angular momenta of unpaired electron couple into a total angular momentum $J=L+S$, which usually turns out to be a good quantum number. When a RE ion is placed in a crystalline environment, the Coulomb interaction with surrounding ligands generates an electrostatic crystal electric field (CEF), breaking the inherent ($2J+1$)-fold degeneracy. Stevens first developed a theoretical approach to describe this phenomenon in 1952 Stevens (1952), introducing angular momentum operators known as Stevens operators to express the ion’s electrostatic potential and directly obtaining the $g$ tensor. This work was later expanded further by Hutchings, who calculated crystal field levels using a simplified point-charge ionic model Hutchings (1964). Consequent studies established that the main characteristic of the CEF Hamiltonian parameters (i.e. Stevens’ operators and their coefficients) is determined solely by the site symmetry of a single RE ion – for example, two or four parameters for cubic, six or eight for hexagonal, and up to twenty-six for lower symmetry environments.
The energy scale of the CEF in rare-earth systems typically spans from a few meV to hundreds of meV, influenced by ligand characteristics and distances from the RE ions. These energy scales often significantly exceed that of exchange interactions, allowing low-energy spin-spin interactions to be treated as perturbations of high-energy single-ion behaviors. Consequently, all relevant exchange tensor components can theoretically be derived using perturbation theory from crystal field wave functions, which, in turn, offers a window into rich physics, spanning the fundamental relations between magnetic moments and crystalline environments, distinct magnetic phenomena for Kramers and non-Kramers magnetic ions, and multipolar properties of the effective magnetic moments. Thus, accurately determining the CEF Hamiltonian and its parameters is the most fundamental step for understanding spin-space anisotropy and overall magnetic characteristics.
In this work, we directly measure the magnetic field $(\mathbf{B})$ dependence of the Zeeman splittings of the CEF levels of CsErSe2 using magneto-topical spectroscopy. From these results, we determine the parameters for the Hamiltonian, i.e. Stevens’ coefficients and the exchange energy scales, to obtain the full field-dependent CEF spectra in two high symmetry directions parallel to the crystalline axes $\hat{b}$ and $\hat{c}$, as shown in Fig. 1 (a) and (b). We also predict the field-induced transition that originates from the crossing of the ground state in $\mathbf{B}\parallel c$ geometry, at $B_{C}\approx 5.4$ T, where magnetization along the $c$-axis displays the field-dependence similar to a metamagnetic transition. Fig. 1(e) shows the calculated magnetization. We verified this transition with the field dependence of AC magnetic susceptibility down to $T=25$ mK, showing that this feature remains robust even below $T_{N}=110$ mK, where long-range ordering sets in at zero field.
Figure 1: CEF energy spectra of single Er3+  ion in CsErSe2 under applied magnetic field (a) $\mathbf{B}\parallel b$ and (b) $\mathbf{B}\parallel c$, calculated from Eq. (1) using the obtained parameters from this study. Eight doublets at zero field ($B=0$) (shaded area on the left side of each plot) are clustered such that five lower levels lie below around 5 meV and the rest above 23 meV. Note that the ground state wave function changes due to the level crossing at $B_{C}=5.4$ T, and another crossing between the upper Zeeman-split level of the ground state and the lower of the first excited occurs at $B=1.4$ T in $\mathbf{B}\parallel c$ shown in (b) (c) Crystal structure of CsErSe2 adopts $P6_{3}/mmc$ space group Xing et al. (2020) (d) Er3+(red) triangular layers formed by edge-sharing ErSe6 octahedra, exhibiting $D_{3d}$ site symmetry (e) The calculated magnetization $M_{b}$ ($\mathbf{B}\parallel b$) and $M_{c}$ ($\mathbf{B}\parallel c$) with mean-field approximation (see the text) at $T=$ 25 mK are shown a function of applied field, which will be discussed in Sec.III.1 and Sec.III.2. The inset displays the same quantities at lower field range with $M_{b}$ and $M_{c}$ calculated at $T=2$ K added. Magnetization curves calculated with the MF approximation at different temperatures are Fig. 3 in the Appendix.
Our work reveals the rich physics of CEF interactions under high magnetic fields, demonstrating that even single-ion physics can host unexpected complexity. This opens a new avenue for exploring the frustrated magnetism unique to rare-earth magnets, where the CEF energy scales are small enough to be manipulated with moderate magnetic field intensity available in a modern laboratory.
The paper is organized as follows. Sec.II describes the CsErSe2 samples, experimental methods, and the Hamiltonian used to calculate energy spectra and magnetic properties. Sec.III.1 presents magneto-optical spectroscopy and analysis for two field directions, from which we determine the Stevens coefficients and spin-exchange energies. Sec.III.2 reports AC susceptibility and magnetization measurements compared with calculations, while Sec.III.3 examines the long-range magnetic order revealed by elastic neutron scattering. Sec.IV discusses the comparison with calculated intensities and the overall implications of our work, and Sec.V summarizes the study. Additional data and analyses are provided in the Appendices.
II Methods
II.1 Materials and Experimental Methods
CsErSe2 crystallizes in a $P6_{3}/mmc$ space group Xing et al. (2020); Scheie et al. (2020) as illustrated in Fig. 1(c). The triangular-lattice planes [Fig. 1(d)] consist of edge-sharing ErSe6 octahedra, allowing $D_{3d}$ site symmetry for Er3+. Within the $ab$-plane, the Se atoms mediate antiferromagnetic (AFM) superexchange interactions between nearest-neighbor Er3+ion pairs, resulting in frustrated triangular magnetism. We identify CsErSe2 indeed exhibits long-range order at zero field at $T_{N}=110$ mK, which we discuss in details in Sec.III.3.
Single crystals of CsErSe2 used in these experiments were grown from powder samples using CsCl flux as described in Xing et al. (2020). Typical sizes of single crystals used for optical spectroscopy are $(1-2)\times(0.3-0.8)\times(0.02-0.04)$ mm3.
[…]
Magneto-infrared spectroscopy was performed in a 17.5 T superconducting magnet with a Fourier-transform infrared (FTIR) spectrometer by Bruker Vertex 80v. Broadband IR light provided by the FTIR spectrometer is guided via evacuated brass light pipes to the sample space, situated at the center of the magnet. Transmitted IR light is delivered to a composite Si-bolometer (IR Labs, Inc.) mounted in the same cryogenic environment, cooled by low-pressure helium gas to a temperature of approximately 5 K. We used both Faraday ($\mathbf{B}\parallel c$) and Voigt ($\mathbf{B}\parallel b$) geometries with the wave vector of the incoming IR light to measure the CsErSe2 transmission spectra. We find that CsErSe2 completely absorbs the spectral range of $130\leq E\leq 210$ cm-1.
AC magnetic susceptibility was measured in a nesting coil set comprised of a primary coil, which generates a very small ac magnetic field, and a pair of sensing coils, coaxially placed within the primary coil. The two sensing coils are wound in opposite directions and connected in series to ensure equal mutual inductance with opposite sign. This results in a very small net induced voltage in a null state. Six single crystal pieces of CsErSe2 samples were stacked together along the $c$-axis and placed in one of the sensing coils. The change of magnetic flux contributed by the sample induces a directly proportional nonzero net voltage. We find the background signal of coils is negligible compared to the signal generated by the presence of the samples. We mount the coil set on a dilution refrigerator such that the DC magnetic field generated by the superconducting magnet is aligned to the $c$-axis.
We measured the neutron diffraction spectra at the HB2A spectrometer at ORNL’s High Flux Isotope Reactor (HFIR) Garlea et al. (2010). We measured 3 g of loose powder of CsErSe2 in a copper can mounted in a dilution refrigerator at temperatures $T=0.5$ K and $T=0.05$ K using $\lambda=2.41$ Å  neutrons. The data are plotted in Fig. 4. The magnetic scattering was isolated by subtracting the high-temperature data from the low-temperature data. We performed Reitveld refinements to the neutron data using the Fullprof software package Rodríguez-Carvajal (1993).
II.2 Single ion Hamiltonian
The single ion Hamiltonian for Er3+under magnetic field is comprised of two terms: the CEF interaction, $\hat{\mathcal{H}}_{\rm CEF}$, which accounts for Coulomb interactions between Er3+and the surrounding anion charge distribution, and Zeeman splitting $\hat{\mathcal{H}}_{\rm Z}$:
$\mathcal{H}=\mathcal{H}_{\rm CEF}+\mathcal{H}_{\rm Z}.$ (1)
In CsErSe2 , the first term can be expressed as a linear combination of the six symmetry-allowed Stevens operators, $\hat{O}^{m}_{n}$, for the $D_{3d}$ site symmetry of the ErSe6 octahedral environment Scheie et al. (2020); Ding et al. (2023)
$\mathcal{H}_{\rm CEF}=B_{2}^{0}\hat{O}_{2}^{0}+B_{4}^{0}\hat{O}_{4}^{0}+B_{4}^{3}\hat{O}_{4}^{3}+B_{6}^{0}\hat{O}_{6}^{0}+B_{6}^{3}\hat{O}_{6}^{3}+B_{6}^{6}\hat{O}_{6}^{6}.$ (2)
At zero field (ZF), the energy eigenstate of this Hamiltonian splits the $J=15/2$ multiplet into eight Kramers doublets, $|n_{\pm}\rangle$ with $n=0,1,2,\cdots,7,8$ [shaded area of Fig. 1(a) and (b)]. Upon applying magnetic field, each doublet splits into two states according to the Zeeman interaction:
$\mathcal{H}_{\rm Z}=-\mu_{B}g_{J}\mathbf{B}\cdot\sum_{i}\hat{\mathbf{J}}_{i},$ (3)
where the Landé $g$-factor, $g_{J}=6/5$, is used for Er3+moments.
Beyond the single-ion picture, spin-spin interactions are present in CsErSe2 . Triangular lattice symmetries of Er3+ ions permit a nearest-neighbor superexchange interaction with XXZ spin symmetry Yamamoto et al. (2014). Since the pseudo-dipolar anisotropic exchange terms Maksimov et al. (2019) vanish within a Weiss mean-field treatment, we confine our analysis to the minimal XXZ Hamiltonian $\hat{\mathcal{H}}_{\rm XXZ}$, featuring nearest-neighbor $J=15/2$ moments coupled by $\mathcal{J}_{\perp}$ and $\mathcal{J}_{z}$ for transverse and longitudinal components:
$\mathcal{H}_{\rm XXZ}=\sum_{\langle ij\rangle}\!\left[\mathcal{J}_{\perp}\!\left(\!\hat{J}_{i,x}\hat{J}_{j,x}+\hat{J}_{i,y}\hat{J}_{j,y}\!\right)\!+\mathcal{J}_{z}\hat{J}_{i,z}\hat{J}_{j,z}\right]\!,$ (4)
in which the indices $\langle i,j\rangle$ refer to nearest-neighbor lattice sites and $\hat{J}_{i,\gamma}$, with $\gamma=x,y,z$, labeling the components of spin $J=15/2$ moments on site $i$ not . The energy scale $\mathcal{H}_{\rm XXZ}$ remains much smaller than the other two single ion contributions in most field and temperature ranges, even below the ordering temperature.
Figure 2: (a,b) Field dependence of normalized FIR absorption intensity at $T=5$ K below 120 cm-1and magnetic field applied in $\mathbf{B}\parallel b$ (a) and $\mathbf{B}\parallel c$ (b). (c) Normalized peak intensity of Raman shift, denoted as $E$, as a function of field in $\mathbf{B}\parallel c$ at $T=5$ K. The FIR spectra in (a)and (b) do not show any field-independent features due to subtracting the ZF spectra. The Raman spectrum, however, displays the field-independent Raman active phonon mode at 49.3 cm-1, which corresponds to the $E_{2g}^{1}$ mode (See Table 2 in Appendix ). The peak positions in (a), (b) and (c) are marked with circles in (d), triangles (e) and crosses in (f). Solid lines in (a-f) indicate transitions from the ground state to field-split excited CEF levels. 1(a-b). Gray-shaded areas mark the regions $B_{\rm ext}>0$ and $E>0$ for clarity.
Finally, we treat the system at finite temperature within a self-consistent Weiss mean-field approximation. Hence, the mean-field Hamiltonian combining Eq. (1) and (4) reduces to a system of $N$ decoupled Er3+ ions as
$\mathcal{H}_{\rm MF}^{\rm sg}$ $=$ $\hat{\mathcal{H}}_{\rm CEF}-{\textstyle\frac{1}{2}}q\left[\mathcal{J}_{\perp}\langle\hat{J}_{y}\rangle^{2}+\mathcal{J}_{z}\langle\hat{J}_{z}\rangle^{2}\right]$ (5)
- $\left[\mu_{B}g_{J}\mathbf{B}-q\left(\mathcal{J}_{\perp}\langle\hat{J}_{y}\rangle\hat{\mathbf{y}}+\mathcal{J}_{z}\langle\hat{J}_{z}\rangle\hat{\mathbf{z}}\right)\right]\!\cdot\!\hat{\mathbf{J}}$,
where $q=6$ is the nearest-neighbor coordination number on the triangular lattice. The magnetization $M_{\alpha}$ is calculated from the mean-field (MF) expectation values of the spin operator $\langle\hat{J}_{\alpha}\rangle$
$M_{\alpha}=g_{J}\mu_{B}\langle\hat{J}_{\alpha}\rangle,$ (6)
This is determined self-consistently via $B=B_{\rm ext}+\lambda M(B)$, where $B_{\rm ext}$ refers to externally applied field and we define the MF parameter $\lambda=\frac{q\mathcal{J}}{g_{J}\mu_{B}}$, where $\mathcal{J}=\mathcal{J}_{\perp}$ or $\mathcal{J}_{z}$, depending on the applied magnetic field orientations Blundell (2001).
III Results
III.1 Magneto-optical spectroscopy
The precise determination Eq. (5) poses a major challenge because it contains eight unknown parameters – six coefficients $\{B_{m}^{n}\}$ of the Stevens operators and two exchange energies, $\mathcal{J}_{\perp}$ and $\mathcal{J}_{z}$. Typically, these parameters have been determined based on inelastic neutron scattering results at zero field. While this provides a direct measure of the CEF levels, there are often more parameters in $\mathcal{H}_{\rm CEF}$ than visible, non-overlapping CEF levels, which leads to a high degeneracy of sets of fitting parameters and hence the uniqueness cannot be guaranteed Li et al. (2017); Bordelon et al. (2020); Zhang et al. (2021); Scheie (2022).
Leveraging the field-dependent Zeeman splittings, we demonstrate a high-fidelity determination of the CEF parameters in CsErSe2 . Fig. 2 displays our main result of magneto-optical spectroscopy, shown below 120 cm-1($\approx$ 15 meV). FIR absorption up to 17.5 T are shown in (a) and (b) for the configurations of $\mathbf{B}\parallel b$ and $\mathbf{B}\parallel c$, respectively. Panel (c) shows the Raman shift under a magnetic field up to 14 T in $\mathbf{B}\parallel c$ geometry. Both IR and Raman data were obtained at the bath temperature at $T=5$ K. For clarity, FIR spectra in (a) and (b) are shown with the ZF spectrum subtracted and normalized to the global maximum, to show only field-dependent CEF levels, while the Raman shift as a function of field, shown in panel (c) is normalized only by the global maximum intensity.
FIR spectra of Fig.2 (a) and (b) display a myriad of crisscrossing energy levels with distinct field dependencies, highly dependent on field orientation, highlighting the strong magneto-anisotropy.
Despite fundamentally different natures of FIR and Raman spectra, the resemblance shown in panels (b) and (c), both taken in $\mathbf{B}\parallel c$, is remarkable, clearly capturing the same physics, except for the field-independent peaks at $E\approx 49$ cm-1in panel (c). It is identified as $E_{2g}^{1}$ phonon from the phonon band calculation (See Table 2 in Appendix).
As mentioned in Sec.II, the FIR spectra of CsErSe2  suffer from complete absorption in $125\leq E\leq 210$ cm-1, above which few discernible field-dependent features were observed. Raman data, however, display discernible features up to $500$ cm-1[ See Fig. 6(a) in Appendix], where transition intensities involving the three highest energy doublets [Fig. 1 (b)] do not appear. This indicates the corresponding intensity is far weaker than those below $E<120$ cm-1, due to exceedingly smaller transition rates. We will return this in Sec. IV.
To obtain the seven parameters at a given field-orientation, we consider only the five lowest doublets, as shown in Fig. 2(d-f). We fit the Raman spectra in panel (f) first with seven parameters – six Stevens’ operators coefficients and $\mathcal{J}_{z}$, and then fit the FIR spectra in $\mathbf{B}\parallel b$ with fixed Stevens Operators’ coefficients to find $\mathcal{J}_{\perp}$. Table 1 lists parameters obtained from the fitting, which agree remarkably well with previous inelastic neutron scattering results Scheie et al. (2020). This is unique to Er3+, a Kramers ion with high $J$ to have seven gaps at ZF that enhance the precision of the fitting Scheie (2021). Using these parameters, we calculated the field evolution of the CEF levels for both field orientations as seen in Fig. 1 (a) and (b). The solid lines shown in Fig. 2 (a-f) are computed by simply subtracting the ground state energy from each excited state energy. This simple procedure alone already captures the peak locations remarkably well, confirming that most features in our spectroscopy data originate from the splitting of the CEF levels.
Table 1: The values of the coefficients of Stevens operators [Eq. (2)] and the exchange energies [Eq. (4)] of CsErSe2  obtained from fits to the spectroscopy data from Fig. 2 (a-c). Units of all shown quantities are $10^{-3}$ meV.
unit [meV] This work Ref. Scheie et al. (2020)
$B_{2}^{0}$ $3.721\times 10^{-2}$ $3.559\times 10^{-2}$
$B_{4}^{0}$ $3.880\times 10^{-4}$ $3.849\times 10^{-4}$
$B_{4}^{3}$ $1.407\times 10^{-2}$ $1.393\times 10^{-2}$
$B_{6}^{0}$ $3.187\times 10^{-6} +$$3.154\times 10^{-6}$
$B_{6}^{3}$ $3.593\times 10^{-6}$ $4.695\times 10^{-6}$
$B_{6}^{6}$ $491\times 10^{-5} +3.3$$381\times 10^{-5}$
$\mathcal{J}_{z}$ $2.64\times 10^{-3}$ $2.4\times 10^{-2}$
$\mathcal{J}_{\perp}$ $0.53\times 10^{-3}$ $0.2\times 10^{-2}$
We now focus on the $\mathbf{B}\parallel c$ spectra, comparing the calculated energy spectra in Fig. 1 (b)] with our data. The most striking feature revealed by calculation is the level crossing on the ground state $B_{C}\approx 5.4$ T, and the crossing of the first excited state at $B_{C}*\approx 1.4$ T. These features are too low in energy for any spectroscopic probe and were not observed in Fig.2(b-c). These level crossings are expected due to the narrowly spaced CEF levels of Er3+even at ZF. For example, at zero field, the gap between ground and the lowest excited state doublets is found to be 0.67 meV, and as the field in $\mathbf{B}\parallel c$, the non-monotonic energy difference between the ground and the first excited state do not exceed this value until the field increases above 12 T.
The direct consequence of the crossing in the ground state is an abrupt change in magnetization, calculated with the MF approximation, as shown in Fig. 1(e): The magnetization along the $c$-axis ($M_{c}$) exhibits a step-like rise at around $B_{C}$, resembling a metamagnetic transition Jacobs and Lawrence (1967). Meanwhile, $M_{b}$, the magnetization along the $b$-axis does not have such a feature. With increasing temperature, the step-like rise in $M_{c}$ disappears quickly, due to increasing thermal populations according to the Boltzmaan factor of the excited states, one of characteristic of the CEF levels.
III.2 Magnetization and AC magnetic susceptibility
As shown in Table 1, the exchange energies in CsErSe2   are at the order $10^{-3}$ meV or less, much smaller than both the CEF interaction and the Zeeman splitting. Consequently, the observed magnetic properties, including conspicuous magnetic anisotropy, are governed by the single-ion ground state properties.
Fig. 3 (a) schematically illustrates the two lowest energy levels under field in $\mathbf{B}\parallel c$, where two field scales of the level-crossing are identified: $B_{C}\approx 5.4$ T and $B_{C}^{*}\approx 1.4$ T, where the ground and the first excited state undergo level-crossing.
In order to examine the corresponding macroscopic magnetic properties, we show the measured AC magnetic susceptibility, $\chi^{\rm ac}_{c}$ (solid lines) in Fig. 3 (b) at various $T$’s between 25 mK and 1 K, with offsets. The calculated $\frac{dM_{c}}{dB}$s are shown in dotted lines, normalized by the $B=0$ values at various $T$’s between 25 mK and 1.008K.
Before we discuss the field dependence of $\chi^{\rm ac}_{c}$, we first note the $\chi^{\rm ac}_{c}(T)$ at ZF in Fig. 3(c), which shows a sharp peak at $T_{C}\approx 110$ mK (red solid line). This is the evidence of long-range magnetic order that is not captured by our single-ion picture. To probe this order, we performed elastic neutron scattering at 50 mK, as discussed in Sec.III.3.
Figure 3: (a) Schematic three lowest energy levels under magnetic field in $\mathbf{B}\parallel c$. Two fields scales at which the level crossing occurs at the 1st excited and the ground states are marked as $B_{C}^{*}$ and $B_{C}$, respectively. (b) Measured AC magnetic susceptibility, $\chi^{\rm ac}_{c}$ (solid lines) are plotted as a function of field in $\mathbf{B}\parallel c$ at various $T$’s as indicated. Dotted lines display $\frac{dM_{c}}{dB}$’s at the corresponding temperature using the calculated $M_{c}$ as shown in Fig. 1(e). Both data and calculations at each $T$ are plotted with a constant offset for clarity. Green and pink dashed lines are guides to eyes for the evolution of the features near $B_{C}^{*}$ and $B_{C}$, respectively. (c) $\chi^{\rm ac}_{c}(T)$ measured at 0.05 T as a function of $T$ is shown in the red solid line, where the spontaneous ordering is identified as a peak at $T=110$ mK. The dashed line shows the $T$ dependence of $\frac{dM_{c}}{dB}$ in $B\rightarrow 0$}, which clearly does not capture the long-range ordering. (d) Calculated $M_{c}$ (dotted lines) and the numerical integration of the $\chi^{\rm ac}_{c}(B)$ (solid line) at $T=25$ mK and 830 mK are plotted as shown. The sharp rise of $M_{c}$ at $B=B_{c}$ is significantly muted from 25 mK and 830 mK in the calculation but the estimated $M_{c}$ from $\chi^{\rm ac}_{c}$ exhibits much less changes between these two $T$’s, where the rise of the $M_{c}$ is significantly stifled, compared to the calculation. They are all normalized by the values at $B=12$ T for comparison. Inset: Experimentally measured magnetization $M_{c}$’s (symbols) are plotted together with the calculated one at 2 (blue), 6 (green), and 10 K (red) in solid lines, which agree well.
Now going back to Fig. 3(b), we find the most prominent feature is a strong local maximum at $B_{C}$ in all observed $T$’s. We identify this as the field-induced transition from one eigenstate to another due to the level crossing. The peak height decreases with increasing $T$, becoming barely discernible at 1.008 K. This level crossing is a property of the ground state, an innate characteristic of the single ion Hamiltonian [Eq. (1)], and hence the peak location is independent of $T$ as marked with a vertical pink potted line.
They are also consistent with a strong peak in the derivatives of the calculated $M_{c}$ ($\frac{dM_{c}}{dB}$), although the feature in the data are considerably muted and broadened compared to $\frac{dM_{c}}{dB}$ at all temperatures. This is even more evident in Fig.3 (d), where the integrated $\chi^{\rm ac}_{c}(B)$ at $T=25$ and 830 mK are plotted in comparison to the calculated $M_{c}$. We believe this is attributed to disorders and defects such that not all Er3+are sitting in the identical $D_{3d}$ octahedral environment, resulting in broadening of the CEF levels as well as their field-dependencies. By $T=$ 2 K, the typical value for the lowest temperature available common magnetometer [inset of Fig.3 (d)], the rapid increase of $M_{c}$ near $B_{C}$ completely vanished.
Another prominent feature in $\chi^{\rm ac}_{c}(B)$ is the rapid decrease of $\chi^{\rm ac}_{c}$ in the vicinity of $B_{C}^{*}$. This behavior smoothly moves to the higher field as $T$ increases, which is in contrast with the unwavering peak location centered at $B_{C}$ in Fig. 3 (b). This behavior is explained as a two-level system with non-monotonic $B$-dependent energy gap, $\Delta_{10}(B)$, which has a local maximum $\Delta_{10}(B_{C}^{*})=0.47$ meV as shown in Fig. 3(a). At a low temperature, this is contrasting to the normal $B$-linear Zeeman splitting in paramagnetic state, where the corresponding magnetization monotonic linear increases, $M=\alpha B$, such that the increase of the magnetization with field exhibits two different slopes in $B<B_{C}^{*}$ and $B>B_{C}^{*}$. This feature is clearly displayed in the calculated magnetization shown in Fig. 3(c) with two distinct slopes in $B<B_{C}$. to decrease as $\Delta_{10}$ increases with field, without In fact, when the system is thermalized at a temperature $T\ll\delta(B_{c}^{*})$, the $\chi^{\rm ac}_{c}$ will approach zero as $B\rightarrow B_{c}^{*}$. As increasing $T$, the thermal populations that undertake this crossing at $B_{C}^{*}$ increase exponentially, and this feature becomes broad and the location of $B_{C}^{*}$ moves higher $B$, shown in the green dashed line.
We notice that there is a sharp dip in $\chi^{\rm ac}_{c}(B)$ in the narrow region of $0<B<B_{C}^{*}$, which was not at all captured by the single-ion Hamiltonian. At this point, we do not have any idea what causes this dip at around $B_{d}=0.2$ T with little $T$-dependence, but it is suspected to be linked with the long-range order, based on its rapid disappearance above $T_{C}$ at 110 mK. Within the ordered phase, the sudden change in magnetic susceptibility may indicate a spin-flop transition Blundell (2001). In the next section, we will further examine the magnetic ordering in CsErSe2   with elastic neutron scattering.
III.3 Magnetic structure analysis with neutron scattering
[…]
The anomaly at $T=110$ mK in $\chi^{\rm ac}_{c}(T)$ [Fig.3(c) ] indicates the formation of long-range order at ZF and prompts us to investigate the nature of the long range ordering in CsErSe2 . In fact, such long-range order has been reported in KErSe2 for powder Xing et al. (2019) and single crystal Ding et al. (2023), and a similar magnetic order is expected. The long-range magnetic order is visible in the neutron scattering as the appearance of new Bragg intensities below $T_{N}$, shown in Fig. 4. These intensities match a $k=(1/2,0,1)$ ordering wavevector, with a very weak $k=(1/2,0,0)$ also present. Refining a magnetic structure on the Er sites to the temperature-subtracted data yielded a coplanar stripe antiferromagnetic order with weak $k=(1/2,0,0)$ correlations, potentially a staggered spin-canting out of plane, shown in Fig. 4 (see Appendix for details). The refined magnetic moment is 3.27(15) $\mu_{B}$, which is only 71(3)% of the maximum static moment from the CEF ground state. Potentially, this is because the magnetic diffraction was taken at 50% of $T_{N}$, and so this moment reduction is likely from a combination of thermal and zero-point fluctuations. Nevertheless, despite the very low transition temperature the magnetic structure is clear. The combination of irreducible representation suggests a weakly first-order transition.
IV discussion
So far, we have shown the anisotropic field evolution of the CEF levels, which enable us to determine the single-ion Hamiltonian with remarkable precision and to make a correct prediction on magnetic behavior. The $g$-factor values are found $g_{\perp}=6.55$ in $\mathbf{B}\parallel b$ and $g_{z}=6.98$, directly calculated from $g_{z(\perp)}=2g_{J}|\langle 0_{\pm}|\hat{J}_{z(\perp)}|0_{\pm(\mp)}\rangle|.$, where $g_{J}$ is Landé $g$-factor of Er3+, $\frac{6}{5}$ and $|0_{\pm}\rangle$ are two doubly degenerate ground states of Eq.(1) $B\rightarrow 0$ limit. Despite the similar $g$-factor values, the magnetic properties are found highly field-orientation-dependent, including the step-like field dependence of the magnetization as a function of field only in $\mathbf{B}\parallel c$ but not in $\mathbf{B}\parallel b$.
Figure 5: (a) Simulated spectrum including selection rules [Eq. (7)] and excitations from thermally populated states (1st and 2nd excited states under field). (b) Transitions with a probability larger than $\rho_{c}=1.7\times 10^{-3}$ (within the spectroscopic resolution) are shown in lines. Different line types show where the excitations are from. The highest-lying three doublets (6 states) with $E>160$ cm-1 have transition probabilities less than 10-4, making it impossible to detect.
Our result reminds us that Zeeman splitting cannot always be assumed to be linear in the applied magnetic field due to two reasons: one is due to a level crossing of the ground state in Fig.3(a), and the other due to a dominant quadratic (second order) component of the Zeeman effect as in the ground state $\mathbf{B}\parallel b$ [Fig. 1(a)] in CsErSe2 . For comparison, the dominant quadratic Zeeman splitting occurs in $\mathbf{B}\parallel c$ for CsYbSe2, a sister compound only with a different RE ionPocs et al. (2021). These two non-linearities make a substantial impact on the $T$ dependence of magnetic susceptibility particularly when $T$ is lower than the first CEF gap at ZF, which is 0.67 meV for CsErSe2 , while 13 meV for CsYbSe2.
One of the most striking features of the field-dependent spectra shown in Fig.2, is multiple crossings and high concentrations of levels within a narrow energy window, i.e., below $E<120$ cm-1. Although the excitations from the ground states alone cannot explain some of the peak positions, narrowly spaced levels imply the thermal population may be responsible for these. In addition, there is an obvious lack of field-dependent spectral lines by three doublets lying above $E>160$ cm-1.
To understand this better, we calculate the transition probability with Fermi’s Golden rule using the single-ion parameters we obtained. This is justified, as the exchange energy among the RE ion is three orders of magnitude smaller than the CEF energy scales, an as a result the eigenstates from the single-ion interactions [Eq. (1)] are responsible for determining the transition probability ($\Gamma_{i\rightarrow f}$), serving as initial($i$) and final ($f$) state:
$\Gamma_{i\rightarrow f}=\frac{2\pi}{\hbar}|\langle f|\mathcal{H}^{p}_{\rm Z}|i\rangle|^{2}\delta(\hbar\omega-\Delta E)p_{\Gamma},$ (7)
where $\mathcal{H}^{p}_{\rm Z}$ refers to the Zeeman interaction between the spin of Er3+and magnetic field by incoming photon, $\hbar\omega$ is the photon energy, $p_{\Gamma}$ is the Bolzmann thermal population factor of the final state replacing the density of state. $\Delta E$ refer to the energy difference between the initial and final states, $\Delta E=E_{f}-E_{i}$. At a given applied magnetic field, Eq. (7) is directly proportional to $|\langle f|\hat{J}_{\alpha}i\rangle|^{2}$, where $\hat{J}_{\alpha}$ denotes the spin component parallel to external magnetic field Furrerer and Podlesnyak (2006); Schaack. (2000)
Fig. 5(a) shows the simulated spectrum at 10K in a colormap to set the probability from Eq. (7) as expected normalized intensity. Overlaid data points are peak locations from FIR and Raman spectroscopy, corresponding to those in Fig.2(e-f), showing that almost all spectra lines are captured in $E<120$ cm-1, provided they have a reasonable transition probability amplitude.
We find that the transition probabilities to the three high-energy levels ($E>160$ cm-1) are at least an order of magnitude smaller $\rho$’ than ones in $E>120$ cm-1, making their detection by optical spectroscopy impossible. Fig. 5(b) displays all allowed transitions for which the normalized transition probability $\rho_{c}>1.7\times 10^{-3}$, where the line type corresponds to initial state. No transition probabilities involving the three high-energy levels were found with this criterion. We also note that discontinuities and abrupt curvature changes of the lines depicted in Fig. 5(b) are the result of the level crossing and consequent abrupt changes of the corresponding eigenstates.
While Raman and far-infrared techniques combined with intense magnetic field have long been used to investigate magnetic couplings and magnonic behavior in $d$-electron systems Xiang et al. (2023); Mou et al. (2024); Lujan et al. (2024); Mai et al. (2025), they prove to be equally valuable for resolving crystal electric field (CEF) excitations in 4f compounds. These methods are particularly effective at revealing magnetoelastic interactions, enhanced by strong spin-orbit coupling, including vibronic bound states formed through coupling between optical phonons and CEF excitations Liu et al. (2018); Čermák et al. (2019). Such interactions originating from the mixing of spin and lattice eigenstates Scheie et al. (2025) give rise to new hybridized quasiparticles with field-dependent spectra and dispersions, which are responsible for, for example, unusual thermal-transport phenomena Pocs et al. (2025).
In CsErSe2 , we did not observe field-dependent coupling between CEF levels and phonon modes, enabling us to determine the single-ion Hamiltonian in a robust manner. Exploring such interactions will help disentangle the interaction with the lattice degrees of freedom at the single-ion level, offering clearer paths to recognize complex spin many-body ground states.
V Summary
We have directly measured the splittings of the doublets CEF levels of CsErSe2 with FIR and Raman spectroscopy under an intense magnetic field, from which we obtain the six Steven’s coefficients and the exchange energies with high precision. We have shown that the magnetic anisotropies of CsErSe2 are prominently manifested in the field-dependent CEF energy spectra: Especially, the ground state under field undergoes level-crossing in $\mathbf{B}\parallel c$ configuration, but not in $\mathbf{B}\parallel b$. This gives rise to a field-induced phase transition in $\mathbf{B}\parallel c$ configuration, where the magnetization abruptly increases at $B_{C}=5.4$ T, manifested as a peak in $\chi^{\rm ac}_{c}(B)$, as the ground state wave function changes across $B_{C}$. Furthermore, the low-lying crossing in the first excited state added the thermally activated drop in $\chi^{\rm ac}_{c}(B)$ at $B_{C}^{*}=1.4$ T, which fades out quickly as increasing $T$.
We also find that CsErSe2 is found to have a spontaneous ordering at $T_{C}=110$ mK in ZF. However, the observed field-orientation dependent magnetic behavior is dictated by a single-ion physics and the spin sector of the Hamiltonian plays a minor or little role in determining macroscopic magnetic properties, even at as low as 25 mK.
Our results show that field-induced level crossings provide a novel route to realizing additional degeneracies, potentially giving rise to unconventional spin ground states. The presence of narrowly spaced CEF levels and their field evolution must also be considered when interpreting field-dependent properties of rare-earth magnets.
Appendix
Figure 6: (a) Raman shift (E) of the full energy spectrum up to 520 cm-1  as a function of field up to 14 T. The region of $E<120$ cm-1  is shown in Fig.  2(c). The purple arrows indicate the field-independent peaks assigned to phonon modes in Table 2. The green arrows mark the field-dependent peaks. Note that the intensities of any field dependent feature lying above 120 cm-1  are significantly smaller than the ones below. (b) The Raman spectra as a function of energy every 2 T step. The data are plotted with a constant offset for the clarity. Purple and green arrows mark the same locations as in panel (b).
Raman Spectra in the extended energy range: Raman shift spectra in extended energy scale (up to 500 cm-1are shown in Fig. 6, under applied field are shown in Fig.6(a) in a colormap of the intensity. In panel (b), the intensities at every 2 T of $B_{\rm ext}$ are plotted as a function of $E$, where each curves are offset by a constant for clarity. The purple arrows $1$ from $8$ show the field-independent peaks that are compared to the modes obtained from the phonon band calculations.
These observed peak locations are listed in Table 2 and they reasonably agrees with the phonon band structure calculation results. The green arrows of $1^{\prime},~2^{\prime}$ and $3^{\prime}$, mark the field-dependent peaks lying above 240 cm-1. We find that none of these peak locations and intensities are matched by the transition probability calculated by Eq. (7) and we do not have a straightforward explanation for these. We speculate that these are from the next $J=13/2$ manifolds or are related to the second order process byproduct from the transitions among other single-ions spin states such as $^{2}H_{11/2}\rightarrow^{4}I_{15/2}$ or $^{4}2S_{3/2}\rightarrow^{4}I_{15/2}$ as their energy(533 nm and 547 nm, respectively) lie very close to the incoming laser energy (532 nm) for the Raman process Taherunnisa et al. (2019).
Table 2: The field-independent peaks indicated with the numbered purple arrows Fig.6 (a) and (b) are listed here and compared to the phonon DFT calculation. The predicted locations’ of the symmetric phonon peaks are shown in the parentheses in the unit of cm-1.
Observed [ cm-1] Assignment (calculated)
1 49.3 $E_{2g}^{1}(46.7)$
2 119.6 unknown
3 125.9 $E_{1g}$ (125.9)
4 132.7 $E_{2g}^{2}$ (127.2)
5 142.8 unknown
6 149.6 $3E_{2g}^{1}$
7 170.3 $A_{1g}$ (164.6)
8 486.3 unknown
not shown 523.4 unknown
Detailed Magnetizations vs Field: Fig. 7 (a) and (b) display the calculated magnetizations of $M_{b}$ and $M_{c}$ at various temperatures as a function of the field intensity in $\mathbf{B}\parallel b$ and $\mathbf{B}\parallel c$ configuration, respectively. Thanks to the ground state crossing in $\mathbf{B}\parallel c$, the low temperature ($T<1$ K) behavior of $M_{c}$ shows the most contrast to that of $M_{b}$, and it mostly washes out by when $T$ reaches 6 K.
Figure 7: Calculated magnetization vs field using the Weiss self-consistent mean field approximation at various temperatures as indicated.
Neutron Refinments: We refined the magnetic structure of CsErSe2 by performing irreducible representational analysis for the $P6_{3}/mmc$ space group and the $(1/2,0,1)$ propagation vector. The four irreducible representations and further information is available in Sup . There are many possible combinations of irreducible representations that will account for this magnetic structure, but we here show the simplest option that preserves uniform moment size: alternating canting along the $c$-axis, shown in Fig. 8.
Figure 8: Refinements of the CsErSe2 magnetic structure with $(1/2,0,1)$ $\Gamma_{1}$ order and with an added $(1/2,0,0)$ canting along the $c$-axis. This additional canting accounts for the $(1/2,0,0)$ Bragg peak, and noticeably lowers the reduced $\chi^{2}$.
Phonon Modes Calculations:All density functional theory (DFT) calculations are performed using the Vienna ab Initio Simulation Package (VASP) version 5.4.4 Kresse and Hafner (1993); Kresse and Furthmüller (1996a, b); Kresse and Hafner (1994) using the projector augmented wave (PAW) Blöchl (1994); Kresse and Joubert (1999) method and the PBE exchange-correlation functional Perdew et al. (1996). A $\Gamma$-centered Monkhorst-Pack Monkhorst and Pack (1976) k-mesh is used for all calculations. Phonons are calculated using the Phonopy software Togo et al. (2023); Togo (2023).
Figure 9: (a) Convergence of phonons with respect to cell size, using a $4\times 4\times 1$ (128 atom) supercell of the primitive cell.
In Fig. 9 we show phonon dispersions for $4\times 4\times 1$ (128 atom) supercells. The close agreement shows that either supercell size could be used, though we use the $4\times 4\times 1$ cell for all final results. Note that longitudinal orbital – transverse orbital (LO-TO) splitting is included in these calculations through the non-analytic term correction (NAC) as implemented in Phonopy. Table 3 listed the $\Gamma$ point energy from Fig. 9 that were used to compare the field-independent Raman shift in Fig. 6. Further information on the calculations and the convergences are available in Sup .
mode # modes Energy (cm-1)
1 N/A 0.00
4-5 $E_{2u}$ 30.30
6-7 $E_{2g}$ 46.72
8-9 $E_{1u}$ 56.41
10 $B_{2u}$ 59.01
11 $B_{1g}$ 82.25
12 $A_{2u}$ 91.34
13-14 $E_{2u}$ 118.15
15-16 $E_{1u}$ 119.25
17-18 $E_{1g}$ 125.91
19-20 $E_{2g}$ 127.17
21 $A_{1g}$ 164.70
22 $A_{2u}$ 168.56
23 $B_{1g}$ 170.46
24 $B_{2u}$ 192.37
Table 3: $\Gamma$ point mode energy in cm-1 as shown in Fig. 9. Extended version of this table is available in Sup .